OpenStudy (hockeychick23):

Pre-calc help please? A skull cleaning factory cleans animal skulls of deer, buffalo, and other types of animals using flesh-eating beetles. The factory owner started with only 7 adult beetles. After 46 days, the beetle population grew to 21 adult beetles. How long did it take before the beetle population was 7,000 beetles?

OpenStudy (photon336):

same kind of problem right exponential growth

OpenStudy (hockeychick23):

ok so a0 is the initial population of 7 a is the population of 7,000 t is the time it takes

OpenStudy (hockeychick23):

\[(\ln(.7*10^3))/1*10^3*x=7000 ? \]

OpenStudy (photon336):

remember we derived this? \[\ln(\frac{ a }{ a_{0} })*\frac{ 1 }{ t } = k\] \[\ln(\frac{ 21 }{ 7 })*(\frac{ 1 }{ 46 }) = 0.02388 = k \]

OpenStudy (photon336):

\[7*e^{0.02388t} = a\]

OpenStudy (photon336):

we know that the final population is 7,000 so we need to solve for that

OpenStudy (hockeychick23):

ok so 7*e^0.02388t=7000

OpenStudy (photon336):

yep, but what I like to do is solve for the variable i'm looking for then plug in \[a_{0}e^{kt} = a \] \[\ln(\frac{ a }{ a_{0} }) = kt \] \[\ln(\frac{ a }{ a_{0} })*\frac{ 1 }{ k } = t \]

OpenStudy (photon336):

\[\ln (\frac{ 7,000 }{ 7 })*(\frac{ 1 }{ 0.02388 }) = t~days \]

OpenStudy (hockeychick23):

ok i got 289 days

OpenStudy (hockeychick23):

thanks!

OpenStudy (photon336):

excellent nice

OpenStudy (photon336):

but a good tip would be to solve the equation for the unknown variable first. then plug in